OpenStudy (anonymous):

if a^2 + a + 1=0 ,then find the value of a^9.

OpenStudy (mathstudent55):

Use the quadratic equation to solve for a, the raise it to the 9th power.

OpenStudy (solomonzelman):

\[ a^2 + a + 1=0\] \[\frac{ -b ± \sqrt{b^2-4(a)(c)} }{ 2a } \] \[\frac{ -1 ± \sqrt{1^2-4(1)(1)} }{ 2 \times 1 } \] \[\frac{ -1 ± i \sqrt{3} }{ 2} \]

OpenStudy (solomonzelman):

now, raise this to the ninth power.